# Efficient Reasoning Workshop

### v1

@@ Abstract
Masked Diffusion Language Models (MDLM) are becoming competitive with their autoregressive counterparts but commonly decode with fixed steps and sequential unmasking. To accelerate decoding, recent works like Fast-dLLM enables parallel decoding via a static global confidence threshold, yet we observe strong block/step-wise confidence fluctuations and, within a dataset, near-identical confidence trajectories across inputs indicated by cosine similarity. Inspired by these two observations, we introduce One-Shot Dynamic Thresholding (OSDT), which calibrates thresholds on a single sequence and applies them to subsequent inputs with negligible overhead. On GPQA, GSM8K, and HumanEval, OSDT attains superior accuracy–throughput trade-offs (+24% tokens/s on GSM8K at the best accuracy, +45% on GPQA with comparable accuracy, and +50% on HumanEval with a modest accuracy gap). Beyond these results, our findings suggest broader opportunities to leverage reusable task-level confidence signatures for more general-purpose algorithmic and systems innovations in diffusion decoding.
@@ 1 Introduction
Masked diffusion language models (MDLM) have recently advanced rapidly, with pre-trained models such as LLaDA-8B (nie2025large, ) and Dream-7B (ye2025dream7b, ) scaling discrete diffusion and demonstrating strong performance across language, math, and code. Inference in these models is semi-autoregressive: the sequence is partitioned into contiguous blocks $B_{1},\dots,B_{T}$ that are generated left-to-right (autoregressive across blocks), while within each block decoding proceeds by diffusion-style denoising steps $s=1,\dots,S$ that can unmask any subset of still-masked positions in a non-left-to-right order. At each step, a mask predictor proposes token distributions and a fixed per‑step quota of token positions is filled with topk by confidence or randomly, while the remaining positions stay masked for later steps.
To push throughput beyond fixed number of tokens per step, recent works like Fast-dLLM (wu2025fastdllm, ) enables parallel decoding by unmasking all tokens above a static global confidence threshold. However, our empirical observations show that confidence is highly non-uniform: it varies across blocks and denoising steps, yet confidence trajectories are remarkably stable across inputs within the same dataset (or, the same task). The first observation calls for adaptivity over the course of generation; the second indicates that a small amount of computation can capture a task-level “confidence signature.” Together, they reveal substantial space for more efficient parallel decoding in MDLM than static schedules can offer.
@@ 4 Evaluation
We evaluate One-Shot Dynamic Thresholding (OSDT) on three benchmarks: GPQA (expert-level Q&A), GSM8K (grade-school math), and HumanEval (code generation). OSDT calibrates a confidence profile on the first sequence of a dataset and applies dynamically adjusted thresholds to subsequent inputs. We study both its hyperparameter behavior and its performance against Fast-dLLM (wu2025fastdllm, ). All experiments use batch size 1 on a single NVIDIA H100 GPU.
@@ 5 Related Work
Recent advances in masked diffusion language models (DLMs) demonstrate their potential as an alternative to autoregressive generation. LLaDA (nie2025large, ) scales discrete diffusion to 8B parameters with block-wise decoding and low-confidence remasking, while Dream (ye2025dream7b, ) improves training via AR initialization and context-adaptive token-level noise rescheduling. Both rely on fixed-step schedules that unmask tokens one by one, limiting efficiency.
Fast-dLLM (wu2025fastdllm, ) addresses this by formally proving when greedy parallel decoding with product-of-marginals is equivalent to sequential decoding in high-confidence regimes, enabling safe parallel generation. It then proposes a confidence-aware parallel decoding rule that unmasks all tokens above a global static cutoff, and introduces prefix and dual (prefix+suffix) KV-Cache designs to improve throughput. However, its thresholding remains task-agnostic and static.

### v2

@@ Abstract
Masked Diffusion Language Models (MDLM) are becoming competitive with their autoregressive counterparts but commonly decode with fixed steps and sequential unmasking. To accelerate decoding, recent works like Fast-dLLM enables parallel decoding via a static global confidence threshold, yet we observe strong block/step-wise confidence fluctuations and, within a dataset, near-identical confidence trajectories across inputs indicated by cosine similarity. Inspired by these two observations, we introduce One-Shot Dynamic Thresholding (OSDT), which calibrates thresholds on a single sequence and applies them to subsequent inputs with negligible overhead. On GPQA, GSM8K, and HumanEval, OSDT attains superior accuracy–throughput trade-offs (+24% tokens/s on GSM8K at the best accuracy, +45% on GPQA with comparable accuracy, and +50% on HumanEval with a modest accuracy gap). Beyond these results, our findings suggest broader opportunities to leverage reusable task-level confidence signatures for more general-purpose algorithmic and systems innovations in diffusion decoding. Our code is available at https://github.com/jackshen-1215/osdt.
@@ 1 Introduction
Masked diffusion language models (MDLM) have recently advanced rapidly, with pre-trained models such as LLaDA-8B (nie2025large) and Dream-7B (ye2025dream7b) scaling discrete diffusion and demonstrating strong performance across language, math, and code. Inference in these models is semi-autoregressive: the sequence is partitioned into contiguous blocks $B_{1},\dots,B_{T}$ that are generated left-to-right (autoregressive across blocks), while within each block decoding proceeds by diffusion-style denoising steps $s=1,\dots,S$ that can unmask any subset of still-masked positions in a non-left-to-right order. At each step, a mask predictor proposes token distributions and a fixed per‑step quota of token positions is filled with topk by confidence or randomly, while the remaining positions stay masked for later steps.
To push throughput beyond fixed number of tokens per step, recent works like Fast-dLLM (wu2025fastdllm) enables parallel decoding by unmasking all tokens above a static global confidence threshold. However, our empirical observations show that confidence is highly non-uniform: it varies across blocks and denoising steps, yet confidence trajectories are remarkably stable across inputs within the same dataset (or, the same task). The first observation calls for adaptivity over the course of generation; the second indicates that a small amount of computation can capture a task-level “confidence signature.” Together, they reveal substantial space for more efficient parallel decoding in MDLM than static schedules can offer.
@@ 4 Evaluation
We evaluate One-Shot Dynamic Thresholding (OSDT) on three benchmarks: GPQA (expert-level Q&A), GSM8K (grade-school math), and HumanEval (code generation). OSDT calibrates a confidence profile on the first sequence of a dataset and applies dynamically adjusted thresholds to subsequent inputs. We study both its hyperparameter behavior and its performance against Fast-dLLM (wu2025fastdllm). All experiments use batch size 1 on a single NVIDIA H100 GPU.
@@ 5 Related Work
Recent advances in masked diffusion language models (DLMs) demonstrate their potential as an alternative to autoregressive generation. LLaDA (nie2025large) scales discrete diffusion to 8B parameters with block-wise decoding and low-confidence remasking, while Dream (ye2025dream7b) improves training via AR initialization and context-adaptive token-level noise rescheduling. Both rely on fixed-step schedules that unmask tokens one by one, limiting efficiency.
Fast-dLLM (wu2025fastdllm) addresses this by formally proving when greedy parallel decoding with product-of-marginals is equivalent to sequential decoding in high-confidence regimes, enabling safe parallel generation. It then proposes a confidence-aware parallel decoding rule that unmasks all tokens above a global static cutoff, and introduces prefix and dual (prefix+suffix) KV-Cache designs to improve throughput. However, its thresholding remains task-agnostic and static.